OpenStudy (anonymous):

Find the velocity vector of a bag just before it hits the ground. It is dropped 500m above the ground and has velocity 5 m/s. The time it takes to hit the ground is 9.6 seconds.

OpenStudy (anonymous):

Assume starts with v0= 5 m/s v = v0 + g t g = 9.8 m/s^2 and t = 9.6 s plug into eqn and note that velocity vector has direction, downward